OpenStudy (anonymous):

Find the point(s) on the graph of y=1/x where the graph is parallel to the line 4x + 9y = 3.

OpenStudy (anonymous):

the slope of you line is \[-\frac{4}{9}\] if i am not mistaken take the derivative, set it equal to \(-\frac{4}{9}\) and solve for \(x\)

OpenStudy (anonymous):

do you know the derivative of \(\frac{1}{x}\) ?